OpenStudy (avanti):

Solve: 2cos^2(theta) -3cos (theta) +1 =0 0≤ theta <2pi please show work, I'm really confused!

OpenStudy (mertsj):

Factor the left side.

OpenStudy (mertsj):

Then set each factor equal to 0 and solve each equation.

OpenStudy (avanti):

so cos(theta) (2cos - 3 ) +1 =0?

OpenStudy (mertsj):

Factor this: \[2x^2-3x+1=0\]

OpenStudy (avanti):

(x-1) (2x-1)

OpenStudy (mertsj):

(x-1) (2x-1)=0

OpenStudy (avanti):

so x =1 and 1/2

OpenStudy (mertsj):

Now set each factor equal to 0 and solve each equation.

OpenStudy (mertsj):

But remember: x is really cos (theta) so replace x with cos (theta) and find theta

OpenStudy (mertsj):

\[\cos \theta =1\]

OpenStudy (mertsj):

\[\cos \theta =\frac{1}{2}\]

OpenStudy (avanti):

how do you find theta.. do you use the unit circle?

OpenStudy (mertsj):

yes

OpenStudy (avanti):

so cosθ=1 is 0 and cosθ=1/2 is √3 /2 ?

OpenStudy (mertsj):

At 0 degrees the x value (which is the cosine) is 1. So theta = 0 degrees At 60 degrees the x value is 1/2 so theta = 60 degrees At 300 degrees the x value is also 1/2 so theta is also 300 degrees.

OpenStudy (avanti):

ohh so those are the three values for x correct?

OpenStudy (avanti):

i just have to convert to rad form right?

OpenStudy (mertsj):

No. They are the values of theta. You are trying to find the angles that have the given cosines.

OpenStudy (mertsj):

yes. 0 degrees is also 0 radians 60 degrees is pi/3 radians 300 degrees is 5pi/3 radians

OpenStudy (avanti):

oh okay i understand it a lot better, thank you :)

OpenStudy (mertsj):

yw